OpenStudy (anonymous):

use the trapezoid rule with n=3 to approximate the area under y=x^2 from x=1 to x=4.

OpenStudy (watchmath):

Divide the interval [1,4] into 3 intervals each of length 1 Then make the table x | 1 2 3 4 y | 1 4 9 16 Then by trapezoid rule (1/2)*1(1+2*2+2*3+4) =(1/2)(1+4+6+4)=15/2.